OpenStudy (anonymous):

3. (i) Differentiate y=x^3-4x^2 Answer: 3x^2-8x (ii) Find the gradient of y= x^3-4x^2 at the point (2,-8). Answer: -4 (iii) Find the equation of the tangent to the curve y= x^3-x^2 at the point (2,-8) Answer: y= -4x (iv) Find the co-ordinates of the other point at which this tangent meets the curve. I need help with this one please! I get stuck when trying to solve he two equations for the tangent and the curve simultaneously.

OpenStudy (phi):

are you asking about part (iv) ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

(iv) Find the co-ordinates of the other point at which this tangent meets the curve. As this graph shows (see below), the tangent meets another point on the curve (but not as a tangent) we have y= -4x y= x^3 -4x^2 and ask where they intersect: -4x = x^3 -4x^2 x^3-4x^2 +4x = 0 x(x^2-4x+4) =0 from which we get x=0 and (x-2)^2=0 i.e. x= 2 thus 0,0 is the other point

OpenStudy (anonymous):

Oh I see now! Thank you very much!